# Distributionally Robust Active Learning for Gaussian Process Regression

### v1

@@ 6.2 Real-World Dataset Experiments
We use the King County house sales [^2] https://www.kaggle.com/datasets/harlfoxem/housesalesprediction, the red wine quality (Cortez and Reis, 2009), and the auto MPG datasets (Quinlan, 1993) (See Appendix C.3 for details). For all experiments, we used SE kernels, where the hyperparameters $\ell$, and $\sigma^{2}$ are adaptively determined by the marginal likelihood maximization (Rasmussen and Williams, 2005) per 10 iterations. The first input is selected uniformly random. Furthermore, we normalize the inputs and outputs of all datasets before the experiments and set $p_{\rm ref}={\cal N}(\boldsymbol{0},0.3\boldsymbol{I}_{d})$.
Figure 2 shows the result. We can confirm the same tendency that DR and CDR variance reductions show superior performance consistently as the synthetic data experiments shown in Figure 1. Note that the fluctuations come from the hyperparameter estimation.
@@ 7 Conclusion
We can consider several future research directions. First, since we do not show the optimality of the convergence rate, developing a (near) optimal algorithm for $E_{T}$ is vital. For this goal, the approximate submodularity (Bian et al., 2017) may be relevant from the empirical superiority of DR variance reduction. Second, since the expectation over $p(\boldsymbol{x}^{*})$ may be intractable, an analysis incorporating the approximation error or developing an algorithm that is efficient without expectation computation may be crucial (DR random does not require the expectation but is often inefficient). Third, although our analyses only require the existence of the maximum over ${\cal P}$, our experiments are limited to the discrete distribution set defined by $L_{\infty}$ ball. Thus, more general experiments regarding, e.g., the continuous probability distributions and the ambiguity sets defined by Kullback-Leibler divergence (Hu and Hong, 2013) and Wasserstein distance (Frogner et al., 2021), are interesting from the practical perspective.
@@ Appendix A Proofs for Section 3
For the first term $\boldsymbol{v}_{t}^{\top}(\cdot)\boldsymbol{f}_{t}$, we follow the proof of Lemma 4 of Vakili et al. [2021a]. Recall the RKHS-based definition of kernel ridge estimator:
@@
for all $\boldsymbol{x},\boldsymbol{z}\in{\cal X}$. Note that since the kernel function, $k$ has partial derivatives due to Assumption 2.8, the derivative of the kernel and the kernel itself are the kernels again as discussed in, e.g., Sec. 9.4 in Rasmussen and Williams [2005] and Sec. 2.2 in Adler [1981]. Thus, we can interpret $\overline{\boldsymbol{v}}_{t}^{\top}(\cdot)\boldsymbol{\epsilon}_{t}$ as the kernel ridge estimator for $g(\boldsymbol{x},1)$, where $\overline{\boldsymbol{v}}_{t}(\boldsymbol{x})=\left(\frac{\partial\boldsymbol{k}_{t}^{\top}(\boldsymbol{x})}{\partial x_{j}}(\boldsymbol{K}_{t}+\sigma^{2}\boldsymbol{I}_{t})^{-1}\right)^{\top}$. In addition, $\|g\|_{{\cal H}_{\overline{k}}}=0$. Therefore, from Theorem 3.11 in Abbasi-Yadkori [2013] and $|g(\boldsymbol{x},1)-\overline{\boldsymbol{v}}_{t}^{\top}(\boldsymbol{x})\boldsymbol{\epsilon}_{t}|=|\overline{\boldsymbol{v}}_{t}^{\top}(\boldsymbol{x})\boldsymbol{\epsilon}_{t}|$, we obtain
@@
That is, the upper bound of the conditional probability given any $(\boldsymbol{x}_{i})_{i\in[t]}$ directly suggests the upper bound of target probability in the left-hand side. Note that from the assumption $(\boldsymbol{x}_{i})_{i\in[t]}$ is independent of $(\epsilon_{i})_{i\in[t]}$ and $f$, the observations $\boldsymbol{y}_{t}$ follows Gaussian distribution even if $(\boldsymbol{x}_{i})_{i\in[t]}$ is fixed.
We leverage Slepian’s inequality shown as Proposition A.2.6 in van der Vaart and Wellner [1996]:
@@
The separability [Definition 5.22 in Handel, 2016] holds commonly. As discussed in Remark 5.23 in Handel [2016], for example, if the sample path is almost surely continuous, then the separability holds. Furthermore, the sample path defined by the commonly used kernel functions, such as linear, SE, and Matérn-$\nu$ kernels with $\nu\geq 1$, is continuous almost surely [Costa et al., 2024]. In addition, if the kernel function is continuous, the posterior mean function is also continuous, almost surely.
@@
Then, we see that the posterior variance of the derivative can be obtained in the same way as the usual GP calculation as follows:
@@
Remained proof is the same as the case of $\mu_{t}^{(j)}$.
@@
As with the existing studies [e.g., Srinivas et al., 2010], we consider the discretization of input space. Let $\overline{{\cal X}}\subset{\cal X}$ be a finite set with each dimension equally divided into $\lceil\tau dr\rceil$, where $\tau>0$. Therefore, $|\overline{{\cal X}}|=\lceil\tau dr\rceil^{d}$ and $\sup_{\boldsymbol{x}\in{\cal X}}\|\boldsymbol{x}-[\boldsymbol{x}]\|_{1}\leq\frac{1}{\tau}$, where $[\boldsymbol{x}]$ is the nearest input in $\overline{{\cal X}}$, that is, $[\boldsymbol{x}]=\operatorname*{arg\,min}_{\tilde{\boldsymbol{x}}\in\overline{{\cal X}}}\|\tilde{\boldsymbol{x}}-\boldsymbol{x}\|_{1}$. Note that we leverage $\overline{{\cal X}}$ purely sake for the analysis, and $\overline{{\cal X}}$ is not related to the algorithm.
@@
As with the existing studies [e.g., Srinivas et al., 2010], we consider the discretization of input space. Let $\overline{{\cal X}}\subset{\cal X}$ be a finite set with each dimension equally divided into $\lceil\tau dr\rceil$, where $\tau>0$. Therefore, $|\overline{{\cal X}}|=\lceil\tau dr\rceil^{d}$ and $\sup_{\boldsymbol{x}\in{\cal X}}\|\boldsymbol{x}-[\boldsymbol{x}]\|_{1}\leq\frac{1}{\tau}$, where $[\boldsymbol{x}]$ is the nearest input in $\overline{{\cal X}}$, that is, $[\boldsymbol{x}]=\operatorname*{arg\,min}_{\tilde{\boldsymbol{x}}\in\overline{{\cal X}}}\|\tilde{\boldsymbol{x}}-\boldsymbol{x}\|_{1}$. Note that we leverage $\overline{{\cal X}}$ purely sake for the analysis, and $\overline{{\cal X}}$ is not related to the algorithm.
@@ C.1 Results for Variance
Figure 3: Result of $\max_{p\in{\cal P}}\mathbb{E}_{p(\boldsymbol{x}^{*})}\left[\sigma_{t}^{2}(\boldsymbol{x}^{*})\right]$ in the synthetic data experiments with $\eta=0,0.001,0.01,0.1$. The horizontal and vertical axes show the number of iterations and $\max_{p\in{\cal P}}\mathbb{E}_{p(\boldsymbol{x}^{*})}\left[\sigma_{t}^{2}(\boldsymbol{x}^{*})\right]$, respectively. The error bar shows mean and standard errors for 20 random trials regarding the random initial point (and the algorithm’s randomness). The top and bottom columns represent the results of the GPR model with SE and Matérn kernels, respectively.
@@ C.3 Details on Real-World Datasets
The King County house sales dataset is a dataset used to predict house prices in King County by 7-dimensional features, such as the area and the number of rooms. This dataset has been used for testing the regression [Park et al., 2020], and a similar dataset has also been used for the AL studies [Park and Kim, 2020]. Although this dataset includes 20000 data, we used randomly chosen 1000 data for simplicity.
Red wine quality dataset [Cortez and Reis, 2009] is a dataset used to predict the quality of wines from the wine ingredients expressed by 11-dimensional features. This dataset includes 1600 data and has been used for the regression problem [Cortez et al., 2009].
Auto MPG dataset [Quinlan, 1993] is a dataset used to predict automobile fuel efficiency from 6-dimensional features, such as the weight of the automobile and engine horsepower. Auto MPG dataset has been used for the AL reserch [Park and Kim, 2020]. This dataset includes 399 data.

### v2

@@ 6.2 Real-World Dataset Experiments
We use the King County house sales [^2] https://www.kaggle.com/datasets/harlfoxem/housesalesprediction, the red wine quality (Cortez and Reis, 2009), and the auto MPG datasets (Quinlan, 1993) (See Appendix C.3 for details). For all experiments, we used SE kernels, where the hyperparameters $\ell$ and $\sigma^{2}$ are adaptively determined by the marginal likelihood maximization (Rasmussen and Williams, 2005) per 10 iterations. The first input is selected uniformly at random. Furthermore, we normalize the inputs and outputs of all datasets before the experiments and set $p_{\rm ref}={\cal N}(\boldsymbol{0},0.3\boldsymbol{I}_{d})$.
Figure 2 shows the result. We can confirm the same tendency that DR and CDR variance reductions show superior performance consistently, as the synthetic data experiments shown in Figure 1. Note that the fluctuations come from the hyperparameter estimation.
@@ 7 Conclusion
We can consider several future research directions. First, since we do not show the optimality of the convergence rate, developing a (near) optimal algorithm for $E_{T}$ is vital. For this goal, the approximate submodularity (Bian et al., 2017) may be relevant from the empirical superiority of DR variance reduction. Second, since the expectation over $p(\boldsymbol{x}^{*})$ may be intractable, an analysis incorporating the approximation error or developing an efficient algorithm without expectation computation may be crucial (DR random does not require the expectation but is often inefficient). Third, although our analyses only require the existence of the maximum over ${\cal P}$, our experiments are limited to the discrete distribution set defined by the $L_{\infty}$ ball. Thus, more general experiments regarding, e.g., the continuous probability distributions and the ambiguity sets defined by Kullback-Leibler divergence (Hu and Hong, 2013) and Wasserstein distance (Frogner et al., 2021), are interesting from the practical perspective.
@@ Appendix A Proofs for Section 3
For the first term $\boldsymbol{v}_{t}^{\top}(\cdot)\boldsymbol{f}_{t}$, we follow the proof of Lemma 4 of [Vakili et al., 2021a]. Recall the RKHS-based definition of kernel ridge estimator:
@@
for all $\boldsymbol{x},\boldsymbol{z}\in{\cal X}$. Note that since the kernel function, $k$ has partial derivatives due to Assumption 2.8, the derivative of the kernel and the kernel itself are the kernels again as discussed in, e.g., Sec. 9.4 in [Rasmussen and Williams, 2005] and Sec. 2.2 in [Adler, 1981]. Thus, we can interpret $\overline{\boldsymbol{v}}_{t}^{\top}(\cdot)\boldsymbol{\epsilon}_{t}$ as the kernel ridge estimator for $g(\boldsymbol{x},1)$, where $\overline{\boldsymbol{v}}_{t}(\boldsymbol{x})=\left(\frac{\partial\boldsymbol{k}_{t}^{\top}(\boldsymbol{x})}{\partial x_{j}}(\boldsymbol{K}_{t}+\sigma^{2}\boldsymbol{I}_{t})^{-1}\right)^{\top}$. In addition, $\|g\|_{{\cal H}_{\overline{k}}}=0$. Therefore, from Theorem 3.11 in [Abbasi-Yadkori, 2013] and $|g(\boldsymbol{x},1)-\overline{\boldsymbol{v}}_{t}^{\top}(\boldsymbol{x})\boldsymbol{\epsilon}_{t}|=|\overline{\boldsymbol{v}}_{t}^{\top}(\boldsymbol{x})\boldsymbol{\epsilon}_{t}|$, we obtain
@@
That is, the upper bound of the conditional probability given any $(\boldsymbol{x}_{i})_{i\in[t]}$ directly suggests the upper bound of the target probability on the left-hand side. Note that from the assumption $(\boldsymbol{x}_{i})_{i\in[t]}$ is independent of $(\epsilon_{i})_{i\in[t]}$ and $f$, the observations $\boldsymbol{y}_{t}$ follows Gaussian distribution even if $(\boldsymbol{x}_{i})_{i\in[t]}$ is fixed.
We leverage Slepian’s inequality shown as Proposition A.2.6 in [van der Vaart and Wellner, 1996]:
@@
The separability [Definition 5.22 in Handel, 2016] holds commonly. As discussed in Remark 5.23 in [Handel, 2016], for example, if the sample path is almost surely continuous, then the separability holds. Furthermore, the sample path defined by the commonly used kernel functions, such as linear, SE, and Matérn-$\nu$ kernels with $\nu\geq 1$, is continuous almost surely [Costa et al., 2024]. In addition, if the kernel function is continuous, the posterior mean function is also continuous, almost surely.
@@
Then, we see that the posterior variance of the derivative can be obtained in the same way as the usual GP calculation, as follows:
@@
The remaining proof is the same as the case of $\mu_{t}^{(j)}$.
@@
As with the existing studies [e.g., Srinivas et al., 2010], we consider the discretization of input space. Let $\overline{{\cal X}}\subset{\cal X}$ be a finite set with each dimension equally divided into $\lceil\tau dr\rceil$, where $\tau>0$. Therefore, $|\overline{{\cal X}}|=\lceil\tau dr\rceil^{d}$ and $\sup_{\boldsymbol{x}\in{\cal X}}\|\boldsymbol{x}-[\boldsymbol{x}]\|_{1}\leq\frac{1}{\tau}$, where $[\boldsymbol{x}]$ is the nearest input in $\overline{{\cal X}}$, that is, $[\boldsymbol{x}]=\operatorname*{arg\,min}_{\tilde{\boldsymbol{x}}\in\overline{{\cal X}}}\|\tilde{\boldsymbol{x}}-\boldsymbol{x}\|_{1}$. Note that we leverage $\overline{{\cal X}}$ purely for the analysis, and $\overline{{\cal X}}$ is not related to the algorithm.
@@
As with the existing studies [e.g., Srinivas et al., 2010], we consider the discretization of input space. Let $\overline{{\cal X}}\subset{\cal X}$ be a finite set with each dimension equally divided into $\lceil\tau dr\rceil$, where $\tau>0$. Therefore, $|\overline{{\cal X}}|=\lceil\tau dr\rceil^{d}$ and $\sup_{\boldsymbol{x}\in{\cal X}}\|\boldsymbol{x}-[\boldsymbol{x}]\|_{1}\leq\frac{1}{\tau}$, where $[\boldsymbol{x}]$ is the nearest input in $\overline{{\cal X}}$, that is, $[\boldsymbol{x}]=\operatorname*{arg\,min}_{\tilde{\boldsymbol{x}}\in\overline{{\cal X}}}\|\tilde{\boldsymbol{x}}-\boldsymbol{x}\|_{1}$. Note that we leverage $\overline{{\cal X}}$ purely for the analysis, and $\overline{{\cal X}}$ is not related to the algorithm.
@@ C.1 Results for Variance
Figure 3: Result of $\max_{p\in{\cal P}}\mathbb{E}_{p(\boldsymbol{x}^{*})}\left[\sigma_{t}^{2}(\boldsymbol{x}^{*})\right]$ in the synthetic data experiments with $\eta=0,0.001,0.01,0.1$. The horizontal and vertical axes show the number of iterations and $\max_{p\in{\cal P}}\mathbb{E}_{p(\boldsymbol{x}^{*})}\left[\sigma_{t}^{2}(\boldsymbol{x}^{*})\right]$, respectively. The error bar shows mean and standard errors for 20 random trials regarding the random initial point (and the algorithm’s randomness). The top and bottom rows represent the results of the GPR model with SE and Matérn kernels, respectively.
@@ C.3 Details on Real-World Datasets
The King County house sales dataset is a dataset used to predict house prices in King County by 7-dimensional features, such as the area and the number of rooms. This dataset has been used for testing the regression [Park et al., 2020], and a similar dataset has also been used for the AL studies [Park and Kim, 2020]. Although this dataset includes 20000 data points, we used a random sample of 1000 data points for simplicity.
Red wine quality dataset [Cortez and Reis, 2009] is a dataset used to predict the quality of wines from the wine ingredients expressed by 11-dimensional features. This dataset includes 1600 data points and has been used for the regression problem [Cortez et al., 2009].
Auto MPG dataset [Quinlan, 1993] is a dataset used to predict automobile fuel efficiency from 6-dimensional features, such as the weight of the automobile and engine horsepower. Auto MPG dataset has been used for the AL research [Park and Kim, 2020]. This dataset includes 399 data points.